# When to use Graphs in RAG: A Comprehensive Analysis for Graph Retrieval-Augmented Generation

### v2

@@ 1 Introduction
Large language models (LLMs), like Claude (Anthropic, 2024) and GPT (OpenAI, 2023) series, have surprised the world with their remarkable capabilities in many real-world tasks, like linguistic comprehension (Brown et al., 2020), question answering (Khashabi et al., 2020), mathematical reasoning (Hong et al., 2025), and content generation (Chowdhery et al., 2023; Hong et al., 2024; Zhang et al., 2024a). Despite the success, LLMs are always criticized for their inability to handle knowledge-intensive tasks and the tendency to generate hallucinations (Huang et al., 2023), especially when faced with questions requiring specialized expertise (Chen et al., 2024b; He et al., 2024; Tan et al., 2024). Retrieval-augmented generation (RAG) (Gao et al., 2023; Lewis et al., 2020) has recently offered a promising approach to adapt LLMs for specific or private domains. Rather than retraining LLMs to incorporate new knowledge and updates (Feng et al., 2025; Fang et al., 2025; Jiang et al., 2025; Wang et al., 2024b), RAG enhances these models by leveraging external knowledge from text corpora. This approach enables LLMs to generate responses by leveraging not only their parametric knowledge but also real-time retrieved domain-specific information, thereby providing more accurate and reliable answers (Chen et al., 2024a; Li et al., 2024).
@@
However, traditional RAG systems often face critical challenges when dealing with large-scale, unstructured domain corpora (Edge et al., 2024; Peng et al., 2024). The textual documents in this corpus, collected from different sources, like research papers, textbooks and technical reports, often vary widely in accuracy and completeness (Guo et al., 2025). The information retrieved by RAG systems can be extensive, complex, and lack clear organization, since domain knowledge is typically scattered across multiple documents without clear hierarchical relationships between different concepts (Sun et al., 2024; Zhang et al., 2024b; Ma et al., 2024). Although RAG systems (Borgeaud et al., 2022; Izacard et al., 2023; Jiang et al., 2023) attempt to manage this complexity by dividing documents into smaller chunks for effective indexing, this approach inadvertently sacrifices crucial contextual information, significantly compromising retrieval accuracy and contextual comprehension for complex reasoning (Han et al., 2024; Zhang et al., 2025; Shengyuan et al., 2024).
@@
To address this, graph retrieval-augmented generation (GraphRAG) (Zhang et al., 2025; Peng et al., 2024; Procko & Ochoa, 2024) has recently emerged as a powerful paradigm that leverages external structured graphs to improve LLMs’ capability on contextual comprehension (Han et al., 2024; Zhang et al., 2025). Early efforts, like Microsoft GraphRAG (Edge et al., 2024) and its variant LazyGraphRAG (Darren Edge, 2024), employ hierarchical community-based search and combine local/global querying for comprehensive responses. Building on this, LightRAG (Guo et al., 2024) improves scalability through dual-level retrieval and graph-enhanced indexing, while GRAG (Hu et al., 2024) introduces a soft pruning technique to mitigate the impact of irrelevant entities in retrieved subgraphs and employs graph-aware prompt tuning to help LLMs interpret topological structure. Further extending these capabilities, StructRAG (Li et al., 2024) tailors data structures to specific tasks by dynamically selecting optimal graph schemas, while KAG (Liang et al., 2024) constructs domain expert knowledge using conceptual semantic reasoning and human-annotated schemas, which significantly reduces noise present in OpenIE systems. These strategies used in GraphRAG models significantly improve retrieval precision and contextual depth, enabling LLMs to address complex, multi-hop queries more effectively.
@@
It is crucial to identify the factors that are currently limiting GraphRAG’s real-world performance. However, quantitatively and fairly assessing the role of graph structures in RAG systems is challenging. Current benchmarks, including HotpotQA (Yang et al., 2018), MultiHopRAG (Tang & Yang, 2024) and UltraDomain (Qian et al., 2024), fail to adequately evaluate the effectiveness of graph structures in RAG systems due to fundamental limitations in both their problem design and corpus composition. ❶ First, existing benchmarks lack granular differentiation in task complexity. Existing benchmarks overemphasize retrieval difficulty, locating scattered facts from corpora, while neglecting reasoning complexity, which involves synthesizing interconnected facts into contextually grounded solutions. As shown in Figure 3, they predominantly focus on narrow task categories, such as simple fact retrieval or linear multi-hop reasoning, without systematically capturing the spectrum of challenges encountered in real-world scenarios (Tang & Yang, 2024). For instance, a typical multi-hop question in existing benchmarks might ask, “Who founded Company Kjaer Weis, and in which city was this person born¿‘ This requires only the extraction of several discrete facts and cannot extend to complex scenarios requiring hierarchical reasoning and contextual synthesis. ❷ Second, corpora in existing RAG benchmarks suffer from inconsistent quality and low information density. Many datasets are built on generic sources like Wikipedia or news articles, which lack domain-specific knowledge or explicit logical connections. While some work, like UltraDomain (Qian et al., 2024), has tried to extract domain-specific corpora from textbooks, they often fail to encode implicit hierarchies for real-world applications. This makes it impossible to assess GraphRAG’s core strengths in leveraging domain hierarchies. For example, a corpus with poorly defined conceptual hierarchies or loosely connected entities cannot meaningfully test whether graph-aware retrieval mechanisms improve multi-hop reasoning or preserve contextual coherence during knowledge acquisition. Additionally, the absence of corpora with varying information densities, ranging from tightly structured domain knowledge to loosely organized real-world documents, further restricts the evaluation of graph structures’ scalability and adaptability.
@@ 2.2 Current RAG Benchmarks
Existing benchmarks, such as HotpotQA (Yang et al., 2018), MultiHopRAG (Tang & Yang, 2024) and UltraDomain (Qian et al., 2024), were primarily designed to evaluate traditional text-centric RAG frameworks. While these benchmarks have advanced the field, they exhibit critical limitations when applied to assessing GraphRAG.
@@ 4.1 Generation Accuracy (Q1)
HippoRAG2 (Gutiérrez et al., 2025) 60.14 31.35 53.38 33.42 64.10 70.84 48.28 49.84 30.95
@@
HippoRAG2 (Gutiérrez et al., 2025) 66.28 36.69 61.98 36.97 63.08 46.13 68.05 58.78 51.54
@@ 4.2 Retrieval Performance (Q2)
HippoRAG2(Gutiérrez et al., 2025) 70.29 79.25 69.77 85.75 82.50 87.82 42.18 79.10
@@
HippoRAG2 (Gutiérrez et al., 2025) 78.70 87.96 77.00 80.94 77.40 86.85 61.12 78.64
@@ 4.3 Graph Complexity (Q3)
As illustrated in Figure 7, HippoRAG2 produces significantly denser graphs, with node and edge counts that far surpass other frameworks. Specifically, on the novel dataset, HippoRAG2 has an average of 2,310 edges and 523 nodes, while on the medical dataset, it has an average of 3,979 edges and 598 nodes. This enhanced graph density improves both information connectivity and coverage, ultimately contributing to superior retrieval and generation capabilities. This observation is consistent with the retrieval performance, which shows that HippoRAG2 achieves higher recall than other baseline models.
@@ 5 Conclusion
To ensure the reproducibility of our research, we have made our dataset, evaluation code, and detailed experimental settings publicly available. We have open-sourced the original data for our benchmark and all the code required to replicate the results presented in this paper. These resources are available in an anonymous repository at https://anonymous.4open.science/r/GraphRAG-Benchmark-CE8D/. Furthermore, we provide a detailed description of our dataset construction process in Appendix C. The specific hyperparameter settings used for all baseline models evaluated in our experiments are also fully documented in Appendix H.2.
@@ A.1 Code and Leaderboard
To promote transparency and reproducibility, we have uploaded our code to GitHub at https://github.com/GraphRAG-Bench/GraphRAG-Benchmark. This repository includes the source code and scripts for evaluation, ensuring that researchers have full access to the resources required to reproduce and extend our work. Besides that, the leaderboard and all other related resources and analyses are collected for the community at https://graphrag-bench.github.io/
@@ Appendix C Benchmark Construction
To address the critical limitations of existing RAG evaluation frameworks, we present a novel benchmark to assess GraphRAG systems through comprehensive task hierarchies and structured knowledge integration. Specifically, our benchmark is constructed through six stages that systematically integrate domain-specific logical hierarchies and contextual dependencies to enable precise control over question difficulty. The example and related prompts are illustrated in Figures 12 and 15.
@@ C.2 Logic Mining
HippoRAG2 (Gutiérrez et al., 2025) 60.14 31.35 53.38 33.42 64.10 70.84 48.28 49.84 30.95
@@
HippoRAG2 (Gutiérrez et al., 2025) 66.28 36.69 61.98 36.97 63.08 46.13 68.05 58.78 51.54
@@
HippoRAG2(Gutiérrez et al., 2025) 70.29 79.25 69.77 85.75 82.50 87.82 42.18 79.10
@@
HippoRAG2 (Gutiérrez et al., 2025) 78.70 87.96 77.00 80.94 77.40 86.85 61.12 78.64
@@ C.6 Refinement
HippoRAG2 (Gutiérrez et al., 2025) 54.79 30.16 50.45 30.30 61.14 68.99 40.52 52.24 32.05
@@
HippoRAG2 (Gutiérrez et al., 2025) 64.50 33.92 64.05 33.02 64.71 48.27 60.77 39.33 36.76
@@ Appendix E Dataset Statistics and Visualization
Benchmark
Avg.
Entities
Avg.
Relations
Prop. of
Non-isolated Entities
Avg.
Degree
Prop.
Degree > 1
Prop.
Degree > 2
Prop.
Degree > 3
Avg.
Component Size
@@ Appendix F Evaluation Metrics
where $\alpha$ is the weight parameter, we set it to 0.5 by default. FC is the Factual correctness and SS is Semantic similarity, they are defined as:
@@ H.1 Implementation Details of Representative RAG Models
HippoRAG2 (Gutiérrez et al., 2025) Knowledge Graph Phrase,Passage Query Embedding Phrase, Chunk Literal Text
@@
All of these GraphRAG methods construct graphs and refine retrieval strategies to boost RAG system performance across various specialized tasks. Due to time limits, we only assess several representative GraphRAG models. We will include more SOTA models, like ArchRAG (Wang et al., 2025b) PIKE-RAG (Wang et al., 2025a) MedRAG (Zhao et al., 2025) PathRAG (Chen et al., 2025) DBCopilot (Wang et al., 2025c) LightPROF (Ao et al., 2025) CG-RAG (Hu et al., 2025).

### v3

@@ 1 Introduction
Large language models (LLMs), like Claude (Anthropic, 2024) and GPT (OpenAI, 2023) series, have surprised the world with their remarkable capabilities in many real-world tasks, like linguistic comprehension (Brown et al., 2020), question answering (Khashabi et al., 2020), mathematical reasoning (Hong et al., 2025), and content generation (Chowdhery et al., 2023; Hong et al., 2024; Zhang et al., 2024a). Despite the success, LLMs are always criticized for their inability to handle knowledge-intensive tasks and the tendency to generate hallucinations (Huang et al., 2023), especially when faced with questions requiring specialized expertise (Chen et al., 2024b; He et al., 2024; Tan et al., 2024). Retrieval-augmented generation (RAG) (Gao et al., 2023; Lewis et al., 2020) has recently offered a promising approach to adapt LLMs for specific or private domains. Rather than retraining LLMs to incorporate new knowledge and updates (Feng et al., 2025; Fang et al., 2025; Jiang et al., 2025; Wang et al., 2024b; Zhang et al., 2025b), RAG enhances these models by leveraging external knowledge from text corpora. This approach enables LLMs to generate responses by leveraging not only their parametric knowledge but also real-time retrieved domain-specific information, thereby providing more accurate and reliable answers (Chen et al., 2024a; Li et al., 2024).
@@
However, traditional RAG systems often face critical challenges when dealing with large-scale, unstructured domain corpora (Edge et al., 2024; Peng et al., 2024). The textual documents in this corpus, collected from different sources, like research papers, textbooks and technical reports, often vary widely in accuracy and completeness (Guo et al., 2025). The information retrieved by RAG systems can be extensive, complex, and lack clear organization, since domain knowledge is typically scattered across multiple documents without clear hierarchical relationships between different concepts (Sun et al., 2024; Zhang et al., 2024b; Ma et al., 2024). Although RAG systems (Borgeaud et al., 2022; Izacard et al., 2023; Jiang et al., 2023) attempt to manage this complexity by dividing documents into smaller chunks for effective indexing, this approach inadvertently sacrifices crucial contextual information, significantly compromising retrieval accuracy and contextual comprehension for complex reasoning (Han et al., 2024; Zhang et al., 2025a; Shengyuan et al., 2024).
@@
To address this, graph retrieval-augmented generation (GraphRAG) (Zhang et al., 2025a; Peng et al., 2024; Procko and Ochoa, 2024) has recently emerged as a powerful paradigm that leverages external structured graphs to improve LLMs’ capability on contextual comprehension (Han et al., 2024; Zhang et al., 2025a). Early efforts, like Microsoft GraphRAG (Edge et al., 2024) and its variant LazyGraphRAG (Darren Edge, 2024), employ hierarchical community-based search and combine local/global querying for comprehensive responses. Building on this, LightRAG (Guo et al., 2024) improves scalability through dual-level retrieval and graph-enhanced indexing, while GRAG (Hu et al., 2024) introduces a soft pruning technique to mitigate the impact of irrelevant entities in retrieved subgraphs and employs graph-aware prompt tuning to help LLMs interpret topological structure. Further extending these capabilities, StructRAG (Li et al., 2024) tailors data structures to specific tasks by dynamically selecting optimal graph schemas, while KAG (Liang et al., 2024) constructs domain expert knowledge using conceptual semantic reasoning and human-annotated schemas, which significantly reduces noise present in OpenIE systems. These strategies used in GraphRAG models significantly improve retrieval precision and contextual depth, enabling LLMs to address complex, multi-hop queries more effectively.
@@
It is crucial to identify the factors that are currently limiting GraphRAG’s real-world performance. However, quantitatively and fairly assessing the role of graph structures in RAG systems is challenging. Current benchmarks, including HotpotQA (Yang et al., 2018), MultiHopRAG (Tang and Yang, 2024) and UltraDomain (Qian et al., 2024), fail to adequately evaluate the effectiveness of graph structures in RAG systems due to fundamental limitations in both their problem design and corpus composition. ❶ First, existing benchmarks lack granular differentiation in task complexity. Existing benchmarks overemphasize retrieval difficulty, locating scattered facts from corpora, while neglecting reasoning complexity, which involves synthesizing interconnected facts into contextually grounded solutions. As shown in Figure 3, they predominantly focus on narrow task categories, such as simple fact retrieval or linear multi-hop reasoning, without systematically capturing the spectrum of challenges encountered in real-world scenarios (Tang and Yang, 2024). For instance, a typical multi-hop question in existing benchmarks might ask, “Who founded Company Kjaer Weis, and in which city was this person born¿‘ This requires only the extraction of several discrete facts and cannot extend to complex scenarios requiring hierarchical reasoning and contextual synthesis. ❷ Second, corpora in existing RAG benchmarks suffer from inconsistent quality and low information density. Many datasets are built on generic sources like Wikipedia or news articles, which lack domain-specific knowledge or explicit logical connections. While some work, like UltraDomain (Qian et al., 2024), has tried to extract domain-specific corpora from textbooks, they often fail to encode implicit hierarchies for real-world applications. This makes it impossible to assess GraphRAG’s core strengths in leveraging domain hierarchies. For example, a corpus with poorly defined conceptual hierarchies or loosely connected entities cannot meaningfully test whether graph-aware retrieval mechanisms improve multi-hop reasoning or preserve contextual coherence during knowledge acquisition. Additionally, the absence of corpora with varying information densities, ranging from tightly structured domain knowledge to loosely organized real-world documents, further restricts the evaluation of graph structures’ scalability and adaptability.
@@ 2.2 Current RAG Benchmarks
Existing benchmarks, such as HotpotQA (Yang et al., 2018), MultiHopRAG (Tang and Yang, 2024) and UltraDomain (Qian et al., 2024), were primarily designed to evaluate traditional text-centric RAG frameworks. While these benchmarks have advanced the field, they exhibit critical limitations when applied to assessing GraphRAG.
@@ 4.1 Generation Accuracy (Q1)
HippoRAG2 (gutiérrez2025HippoRAG2) 60.14 31.35 53.38 33.42 64.10 70.84 48.28 49.84 30.95
@@
HippoRAG2 (gutiérrez2025HippoRAG2) 66.28 36.69 61.98 36.97 63.08 46.13 68.05 58.78 51.54
@@ 4.2 Retrieval Performance (Q2)
HippoRAG2(gutiérrez2025HippoRAG2) 70.29 79.25 69.77 85.75 82.50 87.82 42.18 79.10
@@
HippoRAG2 (gutiérrez2025HippoRAG2) 78.70 87.96 77.00 80.94 77.40 86.85 61.12 78.64
@@ 4.3 Graph Complexity (Q3)
As illustrated in Figure 7, HippoRAG2 produces significantly denser graphs, with node and edge counts that far surpass other frameworks. Specifically, on the novel dataset, HippoRAG2 has an average of 2,310 edges and 523 nodes, while on the medical dataset, it has an average of 3,979 edges and 598 nodes (per 10k corpus tokens). This enhanced graph density improves both information connectivity and coverage, ultimately contributing to superior retrieval and generation capabilities. This observation is consistent with the retrieval performance, which shows that HippoRAG2 achieves higher recall than other baselines.
@@ 5 Conclusion
To ensure the reproducibility of our research, we have made our dataset, evaluation code, and detailed experimental settings publicly available. We have open-sourced the original data for our benchmark and all the code required to replicate the results presented in this paper. These resources are available in the repository at https://github.com/GraphRAG-Bench/GraphRAG-Benchmark. Furthermore, we provide a detailed description of our dataset construction process in Appendix C. The specific hyperparameter settings used for all baseline models evaluated in our experiments are also fully documented in Appendix H.2.
@@ A.1 Code and Leaderboard
To promote transparency and reproducibility, we have uploaded our code to GitHub at https://github.com/GraphRAG-Bench/GraphRAG-Benchmark . This repository includes the source code and scripts for evaluation, ensuring that researchers have full access to the resources required to reproduce and extend our work. In addition, we will continue to maintain and update the repository to reflect future improvements. Besides that, the leaderboard visualization is provided in Figure 10. We have also updated the related resources to the leaderboard.
@@ Appendix C Benchmark Construction
To address the critical limitations of existing RAG evaluation frameworks, we present a novel benchmark to assess GraphRAG systems through comprehensive task hierarchies and structured knowledge integration. Specifically, our benchmark is constructed through six stages that systematically integrate domain-specific logical hierarchies and contextual dependencies to enable precise control over question difficulty.
@@ C.2 Logic Mining
HippoRAG2 (gutiérrez2025HippoRAG2) 60.14 31.35 53.38 33.42 64.10 70.84 48.28 49.84 30.95
@@
HippoRAG2 (gutiérrez2025HippoRAG2) 66.28 36.69 61.98 36.97 63.08 46.13 68.05 58.78 51.54
@@
HippoRAG2(gutiérrez2025HippoRAG2) 70.29 79.25 69.77 85.75 82.50 87.82 42.18 79.10
@@
HippoRAG2 (gutiérrez2025HippoRAG2) 78.70 87.96 77.00 80.94 77.40 86.85 61.12 78.64
@@ C.6 Refinement
HippoRAG2 (gutiérrez2025HippoRAG2) 54.79 30.16 50.45 30.30 61.14 68.99 40.52 52.24 32.05
@@
HippoRAG2 (gutiérrez2025HippoRAG2) 64.50 33.92 64.05 33.02 64.71 48.27 60.77 39.33 36.76
@@ Appendix E Dataset Statistics and Visualization
Benchmark Avg. Entities Avg. Relations Prop. of Non-isolated Entities Avg. Degree Prop. Degree > 1 Prop. Degree > 2 Prop. Degree > 3 Avg. Component Size
@@ Appendix F Evaluation Metrics
where $\alpha$ is the weight parameter, we set it to 0.75 by default. FC is the Factual correctness and SS is Semantic similarity, they are defined as:
@@ G.2 Experiments on open-source model
Table 10: Construction time and all token usage during the indexing phase across different GraphRAG.
Method Time (s) Input Tokens Completion Tokens Total Tokens
MS-GraphRAG (Edge et al., 2024) 292.45 535832 118841 654673
LightRAG (Guo et al., 2024) 710.32 403401 70771 474172
Fast-GraphRAG 281.74 187765 64052 251817
HippoRAG (Gutiérrez et al., 2024) 77.42 154990 22971 177961
HippoRAG2 (gutiérrez2025HippoRAG2) 96.71 283498 46495 329993
KGP (Wang et al., 2024b) 32.01 68540 20675 89215
RAPTOR (Sarthi et al., 2024) 135.21 113641 1900 115541
KET-RAG (Huang et al., 2025) 350.43 433191 84360 517551
LAZY-GraphRAG 253.59 519394 71756 591150
G.3 Experiments on graph construction efficiency
To assess the practical deployment costs of different frameworks, we evaluated the graph construction efficiency on the Novel dataset. We used one book (approximately 56k tokens) as the input corpus and recorded the construction time, total input tokens, and completion tokens during the indexing phase. The quantitative results are presented in Table 10.
We summarize the observations as follows. The efficiency varies significantly across different GraphRAG methods. Notably, KGP demonstrates the highest efficiency, completing the construction process in just 32.01 seconds with the lowest total token usage (89,215). HippoRAG and HippoRAG2 also perform efficiently, maintaining low time and token costs. In contrast, LightRAG and the MS-GraphRAG incur much higher resource consumption. LightRAG requires over 710 seconds to index, while MS-GraphRAG consumes nearly 655,000 tokens. These results suggest that lightweight frameworks offer a far more cost-effective solution for graph indexing, which is critical for applications with limited time or computational budgets.
Table 11: Generation Evaluation (ACC) on Medical dataset across Qwen series models (3B-14B)
Model Fact Retrieval Complex Reasoning Contextual Summarize Creative Generation Avg
GraphRAG (HippoRAG2)
Qwen2.5-3b 60.19 58.11 58.69 51.34 57.08
Qwen2.5-7b 65.65 64.25 64.26 51.85 61.50
Qwen2.5-14b 65.98 62.62 64.95 62.09 63.91
RAG
Qwen2.5-3b 58.66 52.25 58.71 54.89 56.13
Qwen2.5-7b 57.45 53.53 60.52 55.65 56.79
Qwen2.5-14b 57.56 56.01 61.95 60.91 59.10
G.4 Experiments on the impact of LLM size
To investigate the dependence of GraphRAG performance on the underlying LLM capability, we conducted experiments using the Qwen2.5 series (3B, 7B, and 14B) on the Medical dataset. We compared Standard RAG against GraphRAG, employing HippoRAG2 as the representative GraphRAG method. Since the graph index construction is an offline process, we utilized the same pre-constructed index (generated by GPT-4o-mini) used in our main experiments across all model sizes. This experimental design allows us to isolate the impact of model size specifically during the online generation phase. The quantitative results are presented in Table 11.
GraphRAG exhibits a higher sensitivity to model capacity than Standard RAG. While RAG’s performance remains relatively flat across scales (from Avg 56.13 to 59.10), GraphRAG demonstrates substantial growth (from Avg 57.08 to 63.91), indicating a greater reliance on the model’s reasoning ability to synthesize structural information. Notably, we observe a distinct performance inflection point at the 7B parameter scale (Avg increasing from 57.59 to 61.50), suggesting that 7B serves as a practical "minimum size" threshold where the model acquires sufficient reasoning power to effectively leverage graph-based context.
Table 12: Generation Evaluation (ACC) across Novel dataset under different corpus Sizes
Corpus Type Corpus Size Fact Retrieval Complex Reasoning Contextual Summarize Creative Generation
GraphRAG (HippoRAG2)
Small 56k 60.14 53.38 64.10 48.28
Medium 603k 59.99 56.67 65.77 50.06
Large 1132k 59.19 54.29 62.63 51.18
RAG
Small 56k 64.73 58.64 65.75 60.61
Medium 603k 58.43 41.33 62.06 52.54
Large 1132k 58.04 43.20 62.43 47.19
G.5 Experiments on Scalability with Corpus Size
To evaluate the scalability of GraphRAG compared to Standard RAG across varying data volumes, we conducted controlled experiments on the Novel dataset. We defined the corpus sizes based on token counts calculated via tiktoken: Small ($\sim$56k tokens, representing a single book unit), Medium ($\sim$603k tokens), and Large ($\sim$1132k tokens, representing the full dataset). We utilized HippoRAG2 as the representative GraphRAG method and compared it against Standard RAG across these scales. The quantitative results are presented in Table 12.
The comparison reveals a critical insight regarding robustness. Standard RAG suffers significant performance degradation as the corpus size increases. Notably, its accuracy in Complex Reasoning drops heavily from 58.64% (Small) to 43.20% (Large). This supports the hypothesis that vector-based retrieval is prone to capturing high-similarity but irrelevant noise as the search space expands. In contrast, GraphRAG (HippoRAG2) exhibits remarkable stability, maintaining consistent performance across all scales (e.g., Fact Retrieval remains steady at $\sim$60%). This confirms that the performance stability is driven by the structural constraints of the graph (i.e., explicit entity and triple matching), which effectively filter out the retrieval noise that tends to accumulate with increasing corpus size.
@@ H.1 Implementation Details of Representative RAG Models
HippoRAG2 (gutiérrez2025HippoRAG2) Knowledge Graph Phrase,Passage Query Embedding Phrase, Chunk Literal Text
@@
All of these GraphRAG methods construct graphs and refine retrieval strategies to boost RAG system performance across various specialized tasks. Due to time limits, we only assess several representative GraphRAG models. We will include more SOTA models, like ArchRAG (Wang et al., 2025b) PIKE-RAG (Wang et al., 2025a) MedRAG (Zhao et al., 2025) PathRAG (Chen et al., 2025a) DBCopilot (Wang et al., 2025c) LightPROF (Ao et al., 2025) CG-RAG (Hu et al., 2025).